solve
Solve Ý - {{ x(u) do = 2
Answer #1

dL fl4)dn = j- yluldu-L en differnkakig ju-yiui fy= Ge+cie! (Frem O&@) juisget-cet e- s C,= 2+ C, (3) /y4)=6200)€*+cet (From